# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 3 Estimates for Discrete Green’s function
Now, consider the first term of the R.H.S. of (3.9) in the $x$-direction $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.10)
Now, consider the first term of the R.H.S. of (3.9) in the $y$-direction $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.11)
Now, consider the second term of the R.H.S. of (3.9) in the $x$-direction $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\frac{1}{\varepsilon N^{2}}\sqrt{\varepsilon\log(1/\varepsilon)}\;N\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.$ (3.12)
Now, consider the second term of the R.H.S. of (3.9) in the $y$-direction $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;N$
$\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon))^{3}}{N}\leq C\frac{(\log(1/\varepsilon))^{3}}{N}.$ (3.13)
Now, consider the third term of the R.H.S. of (3.9) in the $x$-direction $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\frac{(\log(1/\varepsilon^{3/2}))^{2}}{N^{2}}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\varepsilon^{3/4}\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$ (3.14)
Now, consider the third term of the R.H.S. of (3.9) in the $y$-direction $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{|\Omega_{f,y}|}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\frac{1}{N^{2}}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\;N$
$\leq C\varepsilon^{-3/2}\frac{\log(1/\varepsilon^{3/2})}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}.$ (3.15)
Note that when $(x,y)\in\Omega_{f,y}$, then $x\in[\varepsilon\log(1/\varepsilon),1]$ and $y\in[-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1]$.
If $\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}=e^{-\frac{\alpha x}{\varepsilon}}|_{x=\lambda_{x}}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$ $=e^{-\frac{\alpha}{\varepsilon}\varepsilon\log\left(\frac{1}{\varepsilon}\right)}=e^{-\alpha\log\left(\frac{1}{\varepsilon}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\alpha}}\right)}=\varepsilon^{\alpha}.$ (3.16)
If $\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}=e^{-\frac{\alpha x}{\varepsilon}}|_{x\neq\lambda_{x}}$ and $x_{0}\in(\lambda_{x},1]$, we obtain
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$ $=e^{-\frac{\alpha x_{0}}{\varepsilon}}=\frac{1}{e^{\frac{\alpha x_{0}}{\varepsilon}}}=\frac{1}{1+\frac{\alpha x_{0}}{1!\varepsilon}+\frac{\alpha^{2}x_{0}^{2}}{2!\varepsilon^{2}}+\frac{\alpha^{3}x_{0}^{3}}{3!\varepsilon^{3}}+\dots}\leq n!\varepsilon^{n}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$ $\leq C\varepsilon^{\alpha}.$ (3.17)
Combining (3.15), (3.16) and (3.17), we obtain
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq C\varepsilon^{-3/2}\frac{\log(1/\varepsilon^{3/2})}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}$
$\leq C\varepsilon^{-3/2}\frac{\log(1/\varepsilon^{3/2})}{N}\varepsilon^{\alpha}\leq C\frac{\log(1/\varepsilon^{3/2})}{N}\quad\text{when}\;\alpha\geq 3/2.$ (3.18)
Now, consider the fourth term of the R.H.S. of in the $x$-direction (3.9) $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{c}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{c}}}$
$\leq C\frac{1}{N^{2}}\frac{1}{\sqrt{\varepsilon}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}N\leq C\varepsilon^{-1/2}\frac{1}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}.$ (3.19)
Note that when $(x,y)\in\Omega_{c}$, then $x\in[\varepsilon\log(1/\varepsilon),1]$ and $y\in[-1+\lambda_{y},1-\lambda_{y}]$.
when $x\in[\lambda_{x},1]$
$e^{\frac{-\alpha x}{\varepsilon}}$ $\leq e^{\frac{-\alpha\lambda_{x}}{\varepsilon}}=e^{-\frac{\alpha}{\varepsilon}\varepsilon\log\left(\frac{1}{\varepsilon}\right)}=e^{-\alpha\log\left(\frac{1}{\varepsilon}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\alpha}}\right)}=\varepsilon^{\alpha}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}$ $\leq C\varepsilon^{\alpha}.$ (3.20)
Combining (3.19) and (3.20), we obtain
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\varepsilon^{-1/2}\frac{1}{N}\leq\varepsilon^{\alpha-1/2}\frac{1}{N}\leq\frac{C}{N}.\quad\text{when}\;\alpha\geq 1/2.$ (3.21)
Now, consider the fourth term of the R.H.S. of in the $y$-direction (3.9) $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{c}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{c}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{c}}}$
$\leq C\frac{1}{N^{2}}\frac{1}{\varepsilon^{3/2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}N\leq C\varepsilon^{-3/2}\frac{1}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}.$ (3.22)
Combining (3.22) and (3.20), we obtain
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\frac{\varepsilon^{\alpha-3/2}}{N}\leq\frac{C}{N}.\quad\text{when}\;\alpha\geq 3/2.$ (3.23)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
$\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Lemma 4.3 When $\beta\geq 1/3$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}$.
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})$ $=\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{c}}$ (3.24)
Now, consider the first term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\quad\sqrt{|\Omega_{f,xy}|}\;N\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.25)
Now, consider the first term of the R.H.S. of (3.24) in the $y$- direction $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,xy}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\sqrt{|\Omega_{f,xy}|}\;N\leq C\frac{1}{N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})$
$\leq C\varepsilon^{3/4}\frac{\log(1/\varepsilon^{3/2})}{N}\leq C\frac{\log(1/\varepsilon^{3/2})}{N}.$ (3.26)
Now, consider the second term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;N$
$\leq C\varepsilon^{1/2}\frac{1}{N\varepsilon}\varepsilon^{1/2}(\log(1/\varepsilon)^{1/2}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.$ (3.27)
Now, consider the second term of the R.H.S. of (3.24) in the $y$-direction $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\quad\sqrt{|\Omega_{f,x}|}\;N\leq C\frac{1}{N}\varepsilon^{2}(\log(1/\varepsilon))^{2}\leq C\frac{(\log(1/\varepsilon))^{2}}{N}.$ (3.28)
Now, consider the third term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\quad\sqrt{|\Omega_{f,y}|}\;N\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{N}\varepsilon^{1/4}(\log(1/\varepsilon^{3/2})^{1/2}$
$\leq C\varepsilon^{3/4}\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$ (3.29)
Now, consider the third term of the R.H.S. of (3.24) $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,y}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\sqrt{|\Omega_{f,y}|}\;N\leq C\frac{1}{N}\varepsilon^{1/4}(\log(1/\varepsilon^{3/2})^{1/2}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{N}.$ (3.30)
Now, consider the fourth term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq C\frac{\sqrt{\varepsilon}}{\varepsilon N^{2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;N$
$\leq C\frac{1}{\varepsilon^{1/2}N}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}.$ (3.31)
when $(x,y)\in\Omega_{c}$, then $x\in[\varepsilon\log\left(\frac{1}{\varepsilon}\right),1]$ and $y\in\left[-1+\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right),1-\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)\right]$.
Then say $y=-1+\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})$, we obtain
$1-y$ $=1-\left(-1+\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)\right)=2-\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)$
$1-y$ $\leq 2.\quad\quad\text{and }\quad 1+y\geq 1.$ (3.32)
For sufficiently small $\varepsilon>0$, we obtain
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}$ $\leq e^{-\frac{\beta}{\sqrt{\varepsilon}}}=\frac{1}{e^{\frac{\beta}{\sqrt{\varepsilon}}}}=\frac{1}{1+\frac{\beta}{1!\sqrt{\varepsilon}}+\frac{\beta^{2}}{2!\sqrt{\varepsilon}^{2}}+\frac{\beta^{3}}{3!\sqrt{\varepsilon}^{3}}+\dots}$
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}$ $\leq C\varepsilon^{\eta}\quad\quad\text{for any}\quad\eta>0.$ (3.33)
Also $1+y=\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})$, we obtain
$e^{-\frac{\beta}{\sqrt{\varepsilon}}\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)}$ $\leq e^{-\beta\log\left(\frac{1}{\varepsilon^{3/2}}\right)}\leq e^{\log\left(\frac{1}{\varepsilon^{\frac{-3\beta}{2}}}\right)}\leq e^{\log\left(\varepsilon^{\frac{3\beta}{2}}\right)}\leq\varepsilon^{\frac{3\beta}{2}}.$ (3.34)
Now, combining (3.31),(3.33) and (3.34), we obtain
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\frac{1}{\varepsilon^{1/2}N}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}$
$\leq C\frac{\varepsilon^{-1/2}}{N}\varepsilon^{\frac{3\beta}{2}}\leq\frac{C}{N}.\quad\text{when}\;\beta\geq 1/3.$ (3.35)
Now, consider the fourth term of the R.H.S. of (3.24) in the $y$-direction $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{c}}$

### v2

@@ 4 Estimates related to diffusion terms
Now, estimate the first term of the R.H.S. of (4.18) in the $x$-direction is $|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,xy}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}.$ (4.19)
Now, estimate the first term of the R.H.S. of (4.18) in the $y$- direction is $|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,xy}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,xy}}$
@@
$\leq\frac{C\varepsilon^{1/2}}{N^{2}\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,xy}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{3/4}\frac{(\log(1/\varepsilon^{3/2}))^{3/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3/2}}{N^{3/2}}.$ (4.20)
Now, estimate second term of the R.H.S. of (4.18) in the $x$-direction is $|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.4 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,x}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\frac{\varepsilon^{1/2}}{N^{2}}\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{3/2}\varepsilon}\varepsilon^{1/2}\log(1/\varepsilon)\;\varepsilon^{\frac{3\beta}{2}}\leq C\frac{\log(1/\varepsilon)}{N^{3/2}}.$ (4.21)
Now, estimate second term of the R.H.S. of (4.18) in the $y$-direction is $|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.4 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,x}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\sqrt{|\Omega_{f,x}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{1}{N^{3/2}}\varepsilon^{2}(\log(1/\varepsilon))^{5/2}\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N^{3/2}}.$ (4.22)
Now, estimate the third term of the R.H.S. of (4.18) in the $x$-direction is $|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,y}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{N}\varepsilon^{1/4}(\log(1/\varepsilon^{3/2})^{1/2}\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{3/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}.$ (4.23)
Now, estimate the third term of the R.H.S. of (4.18) in the $y$-direction is $|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,y}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,y}}$
@@
$\leq\frac{C\varepsilon^{1/2}}{N^{2}}\frac{1}{\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{1}{N^{3/2}}\varepsilon^{1/4}\log(1/\varepsilon^{3/2})\leq C\frac{\log(1/\varepsilon^{3/2})}{N^{3/2}}.$ (4.24)
Now, estimate the fourth term of the R.H.S. of (4.18) in the $x$-direction is $|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq C\frac{\sqrt{\varepsilon}}{\varepsilon N^{2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq\frac{C\varepsilon^{\frac{3\beta}{2}-\frac{1}{2}}}{N^{3/2}}(\log(1/\varepsilon))^{1/2}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}\quad\text{when}\;\beta\geq 1/3.$ (4.25)
Now, estimate the fourth term of the R.H.S. of (4.18) in the $y$-direction is $|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,x})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{c}}|$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{c}}$